# Galled Perfect Transfer Networks

### v1

@@ 1 Introduction
We present galled perfect transfer networks (galled PTNs) which are galled tree-based networks with unidirectional edges that explain a set of characters. We then provide polynomial-time algorithms for two problems. In the galled-completion problem, we ask whether it is possible to complete a given tree with transfers to obtain a galled PTN. We also study the galled-compatibility problem in which, given a set of characters, we must decide whether a galled PTN can explain them. During the process, we provide several structural characterizations of characters that can be displayed on a given tree or a network. To improve the reading experience, we provide a sketch for most proofs, and refer to the appendix for detailed proofs.
We also propose a case study of perfect transfer networks on a real dataset consisting of bacterial species and functional characters obtained the Kyoto Encyclopedia of Genes and Genomes (KEGG) database [KEGG]. We show that the conditions required to explain a set of characters with a galled tree are difficult to achieve, as most characters prevent such an explanation, even when taken individually. On the other hand, the case study lets us see that character losses and unresolved species trees can be responsible for this phenomenon. This leads to future questions such as how to model losses and transfers together, and how character-based transfer prediction can be used to resolve trees.
@@ 2 Preliminaries
A tree $T$ is a network with no underlying cycle. We write $u\preceq_{T}v$ if $v$ is on the path from $\rho(T)$ to $u$, in which case $v$ is an ancestor of $u$ and $u$ a descendant of $v$. Note that $v$ is an ancestor and descendant of itself. Two nodes $u$,$v$ are incomparable in $T$ if none descends from the other, i.e. if neither $u\preceq v$ nor $v\preceq u$. The parent of a node $v$ in $T$ is $p_{T}(v)$. For $v\in V(T)$, we use $T(v)$ for the subtree of $T$ rooted at $v$, that is, $T(v)$ contains $v$ and all of its descendants. We may write $L_{T}(v)$ as a shorthand for $L(T(v))$, or just $L(v)$ if $T$ is understood. The set $L(v)$ is called a clade of $T$.
An LGT network (where LGT comes from Lateral, or Horizontal, Gene Transfers[Cardona_2015] is a network $N=(V,E_{S}\cup E_{T})$, where $\{E_{S},E_{T}\}$ is a specified partition of the edge-set of $N$, such that the subgraph $\mathcal{T}_{N}:=(V,E_{S})$ is a tree with the same set of nodes as $N$. The tree $\mathcal{T}_{N}$ is called the support tree of $N$. The edges in $E_{S}$ are called support edges and the edges in $E_{T}$ are called transfer edges. A node that is the endpoint of a transfer edge is called a transfer node. We assume that for each transfer edge $(u,v)\in E_{T}$, the nodes $u$ and $v$ are incomparable in $\mathcal{T}_{N}$. We also assume that transfer nodes have exactly one out-neighbor in $\mathcal{T}_{N}$. The tree obtained from $\mathcal{T}_{N}$ by suppressing its subdivision nodes is called the base tree of $G$ [^1] Suppressing a subdivision node $u$ with parent $p$ and child $v$ consists of removing $u$ and adding an edge from $p$ to $v$. For $v\in V(N)\setminus\{\rho(N)\}$, we use $p_{N}(v)$ to denote the unique in-neighbor of $v$ in the support tree of $N$. For simplicity, an LGT network that is also a galled tree will be called a galled LGT network.
@@ 2.1 Perfect transfer networks
    for each $C\in\mathcal{C}$, there exists a unique node $v\in V$ that reaches every node $w$ satisfying $C\in l(w)$ (single origin).
@@
Figure 1: (a) A tree and characters $C_{1},C_{2},C_{3}$. (b) A PTN with $T$ as base tree that is not a galled tree (two underlying cycles contain the left child of the root). (c) A galled-completion of $T$. Note that in (b) and (c) for every character $C$, there exists a node (in gray) that transmits to all its vertical descendants, and reaches every leaf in $C$.
Figure 1 shows two PTNs for the same set of characters. It does not exhibit the full $\mathcal{C}$-labeling, but a possible origin of each character is annotated on the internal nodes. In [ptn], it is shown that any set of characters can be explained by some PTN. In fact, any tree can become a PTN by adding enough transfer edges. Our goal is to constrain the PTNs to avoid overcomplicated solutions.
@@ 2.2 Properties of galled PTNs
Let $N=(V,E_{S}\cup E_{T})$ be a galled LGT network and let $v\in V(N)$ be a node with two distinct descendants $x$ and $y$ that are transfer nodes (with $v\in\{x,y\}$ being possible). Then one of the transfer edges incident to $x$ or $y$ has both endpoints in the subtree $\mathcal{T}_{N}(v)$.
@@ 2.3 Differences with reconstructions from bipartitions
As we pointed out in the introduction, Warnow et al. [warnow2024statistically] also independently developed galled tree reconstructions algorithms from character data (which are assumed to be SNPs in the paper). Therein, for each character, the species can be in one of two states, and the states are assumed to correspond to a bipartition of a galled tree $N$, where a bipartition is a partition of the leaves that can be obtained by taking a tree displayed by $N$, removing an edge, and taking the two sets of leaves of the resulting connected components (here, “displayed” means a tree that can be obtained by removing one parent edge from each node of in-degree two). They discuss algorithms to reconstruct a galled tree that contains a set of given characters, that is, that contains all the corresponding bipartitions. More specifically, they show that if all the bipartitions of a galled tree $N$ are known (but $N$ is unknown), then $N$ can be reconstructed in polynomial time. They also show that this approach is statistically consistent if characters evolve on a galled tree and may or may not be transmitted on edges labeled as transfer.
@@
We refer to this as the Perfect Phylogenetic Network (PPN) model, as the notion of explaining a character is the same as in [nakhlehtesis]. If we interpret one state as “presence” and the other as “absence”, this is very similar to our formulation, with two notable exceptions. First, we assume an ordering in states from “absence” to “presence”, whereas no such ordering is imposed in the PPN model. More importantly, the notion of displaying a tree does not distinguish between transfer and vertical descent edges. Figure 2 illustrates this difference. Suppose that a given set of characters (i.e., bipartitions) results in the network shown on the left, and consider the character $C$ denoted by the colored square. In the model of [warnow2024statistically], this character is considered as explained, since the displayed tree on the right shows that the character indeed corresponds to a bipartition. On the other hand, this network does not explain the character under the PTN model (see caption). Roughly speaking, displaying a tree allows removing a vertical edge, whereas we would require transmitting characters along that removed vertical edge. See the figure caption for details. We do note that [warnow2024statistically] consider an evolutionary model in which characters evolve down a tree in which transfer edges are explicitly labeled. However, this is only to describe the statistically consistent model, as unlike us the algorithms do not consider this labeling. Let us also mention that [warnow2024statistically] do not consider the tree completion problem.
@@
Figure 2: (a) An example that shows that galled PPNs are not necessarily galled PTNs. The colored square represents a character present at the two rightmost leaves (or in the PPN formulation, there are two states: with or without a square). (b) A tree displayed by the network, obtained by deleting one edge (we left the resulting nodes with a single child, although they could be suppressed). The illustrated character labeling shows that this network explains the character under the PPN model. As for PTNs, one can see from the definition of $F_{C}(N)$ that under the PTN formulation, no internal node of the initial network can contain the character, and thus that the network cannot explain it.
@@ 3 The Galled Completion Problem
For a character $C\in\mathcal{C}$, we will say that a node $v$ is an origin for $C$ if $v$ reaches every leaf in $C$ in $N-F_{C}(N)$. By Lemma 1, our goal is to ensure that every character has an origin. We first show that in any galled-completion of a tree $T$, an origin of a character $C$ must descend from a FA node (or it could be a transfer node added just above it), and it must “give” its character to all other FAs by transferring just above them.
@@
  • either $w\preceq\alpha_{i}$ or $w$ is a transfer node whose child in $\mathcal{T}_{N}$ is $\alpha_{i}$.
@@
By definition, in $T$ the ancestors of FAs are in $F_{C}(N)$ since they have descendants not in $C$. Thus, an origin $w$ in $N$ must be in a FA subtree, or it could be a transfer node added above when adding transfer edges from $T$ to $N$. If there are multiple FAs, $w$ must be able to reach the other ones, and Lemma 2.2 lets us establish that transfer edges are needed to achieve this. ∎
@@
We focus on the first condition. First observe that if $v\in V(T)$, then $v\in V(N)$ as well. Moreover, the parent $p_{v}$ of $v$ in $T$ is either also the parent of $v$ in $\mathcal{T}_{N}$, or the latter is a transfer node created to complete $T$ into $N$. In the latter case, note that at most one transfer node could could be present on the path between $p_{v}$ and $v$ in $\mathcal{T}_{N}$, as otherwise, there would be two consecutive transfer nodes which, as one can check, would violate Lemma 2.2. In other words, some edges of $T$ are subdivided to obtain $\mathcal{T}_{N}$, but each edge can be subdivided at most once.
@@
For the second condition, as we know the origin $w$ descends from some $\alpha_{i}$ or is its parent. Let $\alpha_{j}$ be a FA other than $\alpha_{i}$. Let $l$ be a leaf descending from $\alpha_{j}$ in $\mathcal{T}_{N}$. Since $w$ is an origin, it reaches $l$ in $N-F_{C}(N)$ and thus also in $N$. As $w$ is not an ancestor of $l$, by Lemma 2.2 there is a transfer edge $(u,v)$ in $N$ such that $u\preceq w$ and $l\preceq v$ in $\mathcal{T}_{N}$. The endpoints of this transfer edge must also be in $N-F_{C}(N)$ for $w$ to be able to use it. Next, note that $p_{T}(\alpha_{j})$ has descendants not in $C$. Thus in $N$, $p_{T}(\alpha_{j})$ and its support tree ancestors are in $F_{C}(N)$, and so it follows that $v$ descends from $\alpha_{j}$ or is a transfer node on its parent edge, in which case $v=p_{N}(\alpha_{j})$ (again since edges can only be subdivided once). If $\alpha_{j}$ is a leaf, the latter must hold and we are done. So assume that $\alpha_{j}$ is not a leaf and let $l_{1},l_{2}$ be two leaves descending from $\alpha_{j}$ in $\mathcal{T}_{N}$. Then there are two transfer edges $(u_{1},v_{1})$ and $(u_{2},v_{2})$ that exist in $N-F_{C}(N)$, where $u_{1},u_{2}$ descend from $w$ and $v_{1},v_{2}$ are ancestors of $l_{1}$ and $l_{2}$, respectively. Since these two transfers go out of the $\mathcal{T}_{N}(w)$ subtree, they must in fact be equal by Lemma 2.2. In other words, $w$ reaches every leaf below $\alpha_{j}$ through a single transfer $(u,v)$ such that $v$ is an ancestor of all those leaves. Since $v$ is in $N-F_{C}(N)$, it follows that $v$ must be the parent of $\alpha_{j}$ (and not $\alpha_{j}$ itself because it cannot be a transfer node as its has at least two children, by assumption on $T$). ∎∎
We next argue that the galled requirement places an important limitation on FAs, as there can be at most two per character. This means that a character $C$ must either be a clade of $T$, or it could be split in two clades.
@@
Let $T$ be a galled-completable tree for a character set $\mathcal{C}$. Then $\alpha_{T}(C)\leq 2$ for any character $C\in\mathcal{C}$.
@@
  • Case 2: $a_{2}$ and $b_{2}$ are incomparable. By Lemma 3, there is one transfer with one endpoint being $p_{N}(a_{1})$ or a descendant of $a_{1}$, and the other being $p_{N}(a_{2})$ or a descendant of $a_{2}$. Likewise, there is a similar transfer connecting $b_{1}$ and $b_{2}$. Since $a_{2}$ and $b_{2}$ are incomparable, these two transfers must be distinct, they are both in $\mathcal{T}_{N}(a_{1})$ or $\mathcal{T}_{N}(p_{N}(a_{1}))$, and they both have external endpoints, contradicting Lemma 2.2.
@@
Note that the first statement of the proof of Lemma 3 requires transfer edges to be unidirectional. If we allowed bidirectional transfer edges, then we can devise examples in which the lemma does not hold (not shown due to space constraints). Allowing such edges already leads to more complex structures.
@@
In the forward direction, if $T$ can be completed into some LGT network $N^{\prime}$, we can argue that for every transfer edge $(u,v)$ added to $N$, there is a corresponding transfer edge in $N^{\prime}$ whose endpoints are either $u$ and $v$, or their support tree descendants. Roughly speaking, this means that the cycles of $N$ have corresponding cycles in $N^{\prime}$, and since $N^{\prime}$ is a galled tree, so is $N$. Conversely, if $N$ is galled, it explains $\mathcal{C}$ by construction. That is, for every character split into two clades in $T$, we either added an edge above the subtrees to connect them, or there is an edge that was added in the descendants due to some minimal FA neighbor, which allow meeting the connectivity requirements of Lemma 1. ∎
@@
We first map each transfer edge $(u^{\prime},v^{\prime})$ of $N$ to a distinct transfer edge $(x,y)$ of $N^{\prime}$, such that $y\in\{u^{\prime},v^{\prime}\}$. Let $(u^{\prime},v^{\prime})\in E_{T}$. Note that by the construction of $N$, $u^{\prime},v^{\prime}$ are subdivision nodes of $\mathcal{T}_{N}$ and are parents of FAs $u,v$. Moreover, $u,v$ are FA neighbors, and thus there is some character $C$ with $\alpha_{T}(C)=\{u,v\}$. By Lemma 3, in $N^{\prime}$ there is a transfer edge $(x,y)$, where either $x\preceq_{\mathcal{T}_{N}}u^{\prime}$ and $y=v^{\prime}$, or $x\preceq_{\mathcal{T}_{N}}v^{\prime}$ and $y=u^{\prime}$. Either way, we map $(u^{\prime},v^{\prime})$ to $(x,y)$. Note that by Lemma 3, $u^{\prime}$ and $v^{\prime}$ are only incident to one transfer edge $(u^{\prime},v^{\prime})$. Therefore, only $(u^{\prime},v^{\prime})$ can be mapped to $(x,y)$, as $y\in\{u^{\prime},v^{\prime}\}$.
Next, consider $(u^{\prime},v^{\prime})\in E_{T}$ and its corresponding edge $(x,y)$ in $N^{\prime}$. Suppose that $(x,y)=(\tilde{u},v^{\prime})$ for some descendant $\tilde{u}$ of $u^{\prime}$ in $\mathcal{T}_{N^{\prime}}$. Consider the network $\tilde{N}$ obtained from $N^{\prime}$ by removing $(\tilde{u},v^{\prime})$ and inserting $(u^{\prime},v^{\prime})$ (which does nothing if $\tilde{u}=u^{\prime}$). Note that in $N^{\prime}$, the transfer edge $(\tilde{u},v^{\prime})$ belongs to a unique underlying cycle $H$ formed by the transfer edge, plus the paths from $\tilde{u}$ and from $v^{\prime}$ to the lowest common ancestor of $\tilde{u}$ and $v^{\prime}$ in $\mathcal{T}_{N^{\prime}}$. Since $u^{\prime}$ is on the path between that ancestor and $\tilde{u}$, in $\tilde{N}$ the incorporation of $(u^{\prime},v^{\prime})$ only creates an underlying cycle whose set of vertices is a subset of $H$. Since $H$ did not intersect with other underlying cycles in $N^{\prime}$, this new underlying cycle does not either, and it follows that $\tilde{N}$ is a galled tree as well. The same idea applies if $y=u^{\prime}$ instead.
Since all underlying cycles of $N^{\prime}$ are independent, we can apply the transformation from the previous paragraph to every mapped transfer edge of $N^{\prime}$. Notice that because each $(u^{\prime},v^{\prime})\in E_{T}$ maps to a distinct transfer edge of $N^{\prime}$, no edge of $N^{\prime}$ is “moved” twice by this process. Therefore, we obtain a galled tree whose set of transfer edges contains $E_{T}$. It then follows that $N$ is a galled tree as well.

### v2

@@ 1 Introduction
We present galled perfect transfer networks (galled PTNs) which are galled tree-based networks with unidirectional edges that explain a set of characters. We then provide polynomial-time algorithms for two problems. In the galled-completion problem, we ask whether it is possible to complete a given tree with transfers to obtain a galled PTN. We also study the galled-compatibility problem in which, given a set of characters, we must decide whether a galled PTN can explain them. During the process, we provide several structural characterizations of characters that can be displayed on a given tree or a network. Althought most of our theoretical results are intuitive, the detailed proofs are sometimes involved. To improve the reading experience, we provide a sketch for most proofs, and refer to the appendix for the full details.
We also propose a case study of perfect transfer networks on a real dataset consisting of bacterial species and functional characters obtained from the Kyoto Encyclopedia of Genes and Genomes (KEGG) database [KEGG]. We show that the conditions required to explain a set of characters with a galled tree are difficult to achieve, as most characters prevent such an explanation, even when taken individually. On the other hand, the case study lets us see that character losses and unresolved species trees can be responsible for this phenomenon. This leads to future questions such as how to model losses and transfers together, and how character-based transfer prediction can be used to resolve trees.
@@ 2 Preliminaries
A tree $T$ is a network with no underlying cycle. We write $u\preceq_{T}v$ if $v$ is on the path from $\rho(T)$ to $u$, in which case $v$ is an ancestor of $u$ and $u$ a descendant of $v$ (we may drop the $T$ subscript if unambiguous). Note that $v$ is an ancestor and descendant of itself. Two nodes $u$,$v$ are incomparable in $T$ if none descends from the other, i.e., if neither $u\preceq v$ nor $v\preceq u$. The parent of a node $v$ in $T$ is $p_{T}(v)$. For $v\in V(T)$, we use $T(v)$ for the subtree of $T$ rooted at $v$, that is, $T(v)$ contains $v$ and all of its descendants. We may write $L_{T}(v)$ as a shorthand for $L(T(v))$, or just $L(v)$ if $T$ is understood. The set $L(v)$ is called a clade of $T$.
An LGT network (where LGT comes from Lateral Gene Transfers[Cardona_2015] is a network $N=(V,E_{S}\cup E_{T})$, where $\{E_{S},E_{T}\}$ is a specified partition of the edge-set of $N$, such that the subgraph $\mathcal{T}_{N}:=(V,E_{S})$ is a tree with the same set of nodes as $N$. The tree $\mathcal{T}_{N}$ is called the support tree of $N$. The edges in $E_{S}$ are called support edges and the edges in $E_{T}$ are called transfer edges. A node that is the endpoint of a transfer edge is called a transfer node. We assume that for each transfer edge $(u,v)\in E_{T}$, the nodes $u$ and $v$ are incomparable in $\mathcal{T}_{N}$. We also assume that transfer nodes have exactly one out-neighbor in $\mathcal{T}_{N}$. The tree obtained from $\mathcal{T}_{N}$ by suppressing its subdivision nodes is called the base tree of $N$ [^1] Suppressing a subdivision node $u$ with parent $p$ and child $v$ consists of removing $u$ and adding an edge from $p$ to $v$. For $v\in V(N)\setminus\{\rho(N)\}$, we use $p_{N}(v)$ to denote the unique in-neighbor of $v$ in the support tree of $N$. For simplicity, an LGT network that is also a galled tree will be called a galled LGT network.
@@ 2.1 Perfect transfer networks
    for each $C\in\mathcal{C}$, there exists a unique node $v\in V$ that, in $N$, reaches every node $w$ satisfying $C\in l(w)$ (single origin).
@@
Figure 1: (a) A tree and characters $C_{1},C_{2},C_{3}$. (b) A PTN with $T$ as base tree that is not a galled tree (two underlying cycles contain the left child of the root). Notice for example the circle character $C_{2}$, which does not use any transfer and only needs to be transmitted vertically to its descendants after emergence. On the other hand, the square character $C_{1}$ needs to be transferred to two other ancestral species (left and right) after emergence. The triangle character $C_{3}$ needs one transfer, and uses the same arc going to the left as $C_{1}$. (c) A different PTN with $T$ as a base tree, which is a galled-completion of $T$.
In short, each character $C\in\mathcal{C}$ must emerge at some node $v$, then be transmitted to every vertical descendant, and possibly horizontally to other species (and if so, such other species must in turn also transmit to their vertical descendants, and possibly transfer horizontally as well). Figure 1 shows two PTNs for the same set of characters. It does not exhibit the full $\mathcal{C}$-labeling, but a possible origin of each character is annotated on the internal nodes. In [ptn], it is shown that any set of characters can be explained by some PTN. In fact, any tree can become a PTN by adding enough transfer edges. Our goal is to constrain the PTNs to avoid overcomplicated solutions.
@@ 2.2 Properties of galled PTNs
Let $N=(V,E_{S}\cup E_{T})$ be a galled LGT network and let $v\in V(N)$ be a node with two distinct descendants $x$ and $y$ in $\mathcal{T}_{N}$ that are transfer nodes (with $v\in\{x,y\}$ being possible). Then one of the transfer edges incident to $x$ or $y$ has both endpoints in the subtree $\mathcal{T}_{N}(v)$.
@@ 2.3 Differences with reconstructions from bipartitions
As we pointed out in the introduction, Warnow et al. [warnow2024statistically] also independently developed galled tree reconstructions algorithms from character data (which are assumed to be SNPs in the paper). Therein, for each character, the species can be in one of two states, and the states are assumed to correspond to a bipartition of a galled tree $N$. A bipartition is a partition of the leaves that can be obtained by taking a tree displayed by $N$, removing an edge, and taking the two sets of leaves of the resulting connected components (here, “displayed” means a tree that can be obtained by removing one parent edge from each node of in-degree two). They discuss algorithms to reconstruct a galled tree that contains a set of given characters, that is, that contains all the corresponding bipartitions. More specifically, they show that if all the bipartitions of a galled tree $N$ are known (but $N$ is unknown), then $N$ can be reconstructed in polynomial time. They also show that this approach is statistically consistent if characters evolve on a galled tree and may or may not be transmitted on edges labeled as transfer.
@@
We refer to this as the Perfect Phylogenetic Network (PPN) model, as the notion of explaining a character is the same as in [nakhlehtesis]. If we interpret one state as “presence” and the other as “absence”, this is very similar to our formulation, with two notable exceptions. First, in our work on PTNs we impose an ordering in states from “absence” to “presence”, whereas no such ordering is assumed in the PPN model. More importantly, the notion of displaying a tree does not distinguish between transfer and vertical descent edges. Figure 2 illustrates this difference. Suppose that a given set of characters (i.e., bipartitions) results in the network shown on the left, and consider the character $C$ denoted by the colored square. In the model of [warnow2024statistically], this character is considered as explained, since the displayed tree on the right shows that the character indeed corresponds to a bipartition. On the other hand, this network does not explain the character under the PTN model. Roughly speaking, displaying a tree allows removing a vertical edge, whereas we would require transmitting characters along that removed vertical edge. See the figure caption for details. We do note that [warnow2024statistically] consider an evolutionary model in which characters evolve down a tree in which transfer edges are explicitly labeled. However, this is only to describe the statistically consistent model, as unlike us the algorithms do not consider this labeling. Let us also mention that [warnow2024statistically] do not consider the tree completion problem.
@@
Figure 2: (a) An example that shows that galled PPNs are not necessarily galled PTNs. The colored square represents a character present at the two rightmost leaves (or in the PPN formulation, there are two states: with or without a square). (b) A tree displayed by the network, obtained by deleting one edge (we left the resulting subdivision node, although they could be suppressed). The illustrated character labeling shows that this network explains the character under the PPN model. As for PTNs, one can see from the definition of $F_{C}(N)$ that under the PTN formulation, no internal node of the initial network can contain the character, that there is no possible origin, and thus that the network cannot explain it.
@@ 3 The Galled Completion Problem
For a character $C\in\mathcal{C}$ and a network $N$, we will say that a node $v$ is an origin for $C$ if $v$ reaches every leaf in $C$ in $N-F_{C}(N)$. By Lemma 1, our goal is to add arcs to $T$ to ensure that every character has an origin. We first show that in any galled-completion of a tree $T$, an origin of a character $C$ must descend from a FA node (or it could be a transfer node added just above it), and it must “give” its character to all other FAs by transferring just above them.
@@
  • either $w\preceq_{\mathcal{T}_{N}}\alpha_{i}$ or $w$ is a transfer node whose child in $\mathcal{T}_{N}$ is $\alpha_{i}$.
@@
By definition, in $T$ the ancestors of FAs are in $F_{C}(N)$ since they have descendants not in $C$. Thus, an origin $w$ in $N$ must be in a FA subtree, or it could be a transfer node inserted above such an FA during the process of adding transfer edges from $T$ to $N$. If there are multiple FAs, $w$ must be able to reach the other ones, and Lemma 2.2 lets us establish that transfer edges as described are needed to achieve this. ∎
@@
We focus on the first condition. First observe that if $v\in V(T)$, then $v\in V(N)$ as well. Moreover, the the parent of $v$ in $\mathcal{T}_{N}$ is either also the parent $p_{v}$ of $v$ in $T$, or it could be a transfer node created to complete $T$ into $N$. In the latter case, note that at most one transfer node could could be present on the path between $p_{v}$ and $v$ in $\mathcal{T}_{N}$, as otherwise, there would be two consecutive transfer nodes which, as one can check, would violate Lemma 2.2 ($p_{v}$ would have two descendants with outgoing transfers). This lets us establish that some edges of $T$ are subdivided to obtain $\mathcal{T}_{N}$, but each edge can be subdivided at most once.
@@
For the second condition, as we now know the origin $w$ descends from some $\alpha_{i}$ or is its parent. Let $\alpha_{j}$ be a FA other than $\alpha_{i}$. Let $l$ be a leaf descending from $\alpha_{j}$ in $\mathcal{T}_{N}$. Since $w$ is an origin, it reaches $l$ in $N-F_{C}(N)$ and thus also in $N$. Since first-appearance nodes are incomparable in $T$ and $\mathcal{T}_{N}$, $w$ is not an ancestor of $l$ in $\mathcal{T}_{N}$. By Lemma 2.2 there is a transfer edge $(u,v)$ in $N$ such that $u\preceq w$ and $l\preceq v$ in $\mathcal{T}_{N}$. The endpoints of this transfer edge must also be in $N-F_{C}(N)$ for $w$ to be able to use it. Next, note that $p_{T}(\alpha_{j})$ has descendants not in $C$. Thus in $N$, $p_{T}(\alpha_{j})$ and its support tree ancestors are in $F_{C}(N)$, and so it follows that $v$ descends from $\alpha_{j}$ or is a transfer node on its parent edge, in which case $v=p_{N}(\alpha_{j})$ (again since edges can only be subdivided once). If $\alpha_{j}$ is a leaf, the latter must hold and we are done. So assume that $\alpha_{j}$ is not a leaf and let $l_{1},l_{2}$ be two leaves descending from $\alpha_{j}$ in $\mathcal{T}_{N}$. Then there are two transfer edges $(u_{1},v_{1})$ and $(u_{2},v_{2})$ that exist in $N-F_{C}(N)$, where $u_{1},u_{2}$ descend from $w$ and $v_{1},v_{2}$ are ancestors of $l_{1}$ and $l_{2}$, respectively. Since these two transfers go out of the $\mathcal{T}_{N}(w)$ subtree, they must in fact be equal by Lemma 2.2. In other words, $w$ reaches every leaf below $\alpha_{j}$ through a single transfer $(u,v)$ such that $v$ is an ancestor of all those leaves. Since $v$ is in $N-F_{C}(N)$, it follows that $v$ must be the parent of $\alpha_{j}$ (and not $\alpha_{j}$ itself because it cannot be a transfer node as its has at least two children, by assumption on $T$). ∎∎
We next argue that the galled requirement places an important limitation on FAs, as there can be at most two per character. This means that a character $C$ must either be a clade of $T$, or it could be split in two clades, but not more. This is both useful in theory, but also in our experiments since it allows checking quickly whether an individual character can be explained by a galled-tree.
@@
Let $T$ be a galled-completable tree for a character set $\mathcal{C}$. Then for any character $C\in\mathcal{C}$, ${\alpha_{T}(C)\leq 2}$.
@@
  • Case 2: $a_{2}$ and $b_{2}$ are incomparable. By Lemma 3, there is one transfer with one endpoint being $p_{N}(a_{1})$ or a descendant of $a_{1}$, and the other being $p_{N}(a_{2})$ or a descendant of $a_{2}$. Likewise, there is a similar transfer connecting $b_{1}$ and $b_{2}$. Since $a_{2}$ and $b_{2}$ are incomparable, these two transfers must be distinct, they are both in $\mathcal{T}_{N}(a_{1})$ or $\mathcal{T}_{N}(p_{N}(a_{1}))$, and they both have external endpoints since $b_{2}$ is not comparable with $a_{1}$, contradicting Lemma 2.2.
@@
Note that in the detailed proof of the first statement of Lemma 3, we use the fact that transfer edges are unidirectional — the statement in the sketch that $b_{1}$ needs to use the “same” transfer as $a_{1}$ remains true, but they do not need to send the character to the same clade in a bidirectional setting, as the sending could go both ways. If we allowed bidirectional transfer edges, we can devise examples in which the lemma does not hold. This shows that even apparently minor changes to the model can lead to more complex structures.
@@
In the forward direction, if $T$ can be completed into a galled tree $N^{\prime}$ that explains $\mathcal{C}$, we can argue that for every transfer edge $(u,v)$ the is present in the redundancy-free network $N$, there is a corresponding transfer edge in $N^{\prime}$ whose endpoints are either $u$ and $v$, or in their support tree descendants. Roughly speaking, this means that the cycles of $N$ have corresponding cycles in $N^{\prime}$, and since $N^{\prime}$ is a galled tree, so is $N$. Conversely, if $N$ is galled, it explains $\mathcal{C}$ by construction. That is, for every character split into two clades in $T$, we either added an edge above the subtrees to connect them, or there is an edge that was added in the descendants due to some minimal FA neighbor, which allow meeting the connectivity requirements of Lemma 1. ∎
@@
We first build an injective map, that associates each transfer edge $(u^{\prime},v^{\prime})$ of $N$ with a distinct transfer edge $(x,y)$ of $N^{\prime}$, such that $y\in\{u^{\prime},v^{\prime}\}$. Let $(u^{\prime},v^{\prime})\in E_{T}$. Note that by the construction of $N$, $u^{\prime},v^{\prime}$ are subdivision nodes of $\mathcal{T}_{N}$ and are parents of FAs $u,v$. Moreover, $u,v$ are FA neighbors, and thus there is some character $C$ with $\alpha_{T}(C)=\{u,v\}$. By Lemma 3, in $N^{\prime}$ there is a transfer edge $(x,y)$, where either $x\preceq_{\mathcal{T}_{N}}u^{\prime}$ and $y=v^{\prime}$, or $x\preceq_{\mathcal{T}_{N}}v^{\prime}$ and $y=u^{\prime}$. Either way, we map $(u^{\prime},v^{\prime})$ to $(x,y)$. We claim that no other transfer edge of $N$ can be mapped to $(x,y)$ in this manner. Suppose that some other transfer edge $(u^{\prime\prime},v^{\prime\prime})$ of $N$ maps to $(x,y)$. According to our map, $y=u^{\prime\prime}$ or $y=v^{\prime\prime}$, but either way, $y$ is incident to two transfer arcs in $N$, contradicting Lemma 3. Therefore, only $(u^{\prime},v^{\prime})$ can be mapped to $(x,y)$.
Next, consider $(u^{\prime},v^{\prime})\in E_{T}$ and its associated edge $(x,y)$ in $N^{\prime}$. Suppose that $(x,y)=(\tilde{u},v^{\prime})$ for some descendant $\tilde{u}$ of $u^{\prime}$ in $\mathcal{T}_{N^{\prime}}$. Consider the network $\tilde{N}$ obtained from $N^{\prime}$ by removing $(\tilde{u},v^{\prime})$ and inserting $(u^{\prime},v^{\prime})$ (which does nothing if $\tilde{u}=u^{\prime}$). Note that in $N^{\prime}$, the transfer edge $(\tilde{u},v^{\prime})$ belongs to a unique underlying cycle $H$ formed by the transfer edge, plus the paths from $\tilde{u}$ and from $v^{\prime}$ to the lowest common ancestor of $\tilde{u}$ and $v^{\prime}$ in $\mathcal{T}_{N^{\prime}}$. Since $u^{\prime}$ is on the path between that ancestor and $\tilde{u}$, in $\tilde{N}$ the incorporation of $(u^{\prime},v^{\prime})$ only creates an underlying cycle whose set of vertices is a subset of $H$. Since $H$ did not intersect with other underlying cycles in $N^{\prime}$, this new underlying cycle does not either, and it follows that $\tilde{N}$ is a galled tree as well. The same idea applies if $y=u^{\prime}$ instead.
Since all underlying cycles of $N^{\prime}$ are independent, we can apply the transformation from the previous paragraph to replace every mapped transfer edge of $N^{\prime}$ by a transfer edge of $N$, one after another, while maintaining the property that no cycles intersect. Notice that because each $(u^{\prime},v^{\prime})\in E_{T}$ maps to a distinct transfer edge of $N^{\prime}$, every edge of $N$ is incorporated, and no edge of $N^{\prime}$ is “moved” twice by this process. Therefore, we obtain a galled tree whose set of transfer edges contains $E_{T}$. It then follows that $N$ is a galled tree as well.